How do I simplify (5+2 times the square root of 6) times (8-3 times the square root of 6)?
Answer by stanbon(48535)

How do I simplify (5+2 times the square root of 6) times (8-3 times the square root of 6)?
------------------------
(5+2sqrt(6))(8-3sqrt(6))
Use FOIL to get:
F: 5*8 = 40
O: 5*-3sqrt(6)= -15sqrt(6)
I: 8*2sqrt(6) = 16sqrt(6)
L: 2sqrt(6)*-3sqrt6 = -6*6 = - 36
--------------------------
Combine to get:
= 4+sqrt(6)
==================
Cheers,
Stan H.